OpenStudy (anonymous):

Find the slope of the line that is parallel and perpendicular to each of the following equation. EQUATION: 2x – 2y = -6 WORK: PARALLEL SLOPE: PERPENDICULAR SLOPE:

OpenStudy (luigi0210):

Set it equal to y

OpenStudy (anonymous):

Ok so I have a table full of these problems, and just need smeone to show me how to do one and I got the rest !

OpenStudy (anonymous):

ok -2y=-2x-6

OpenStudy (anonymous):

-y=-x-3

OpenStudy (anonymous):

y=x+3

OpenStudy (luigi0210):

@texaschic101 you're a great teacher, why don't you take this one

OpenStudy (texaschic101):

you got it ....thanks

OpenStudy (luigi0210):

Uhm, @aprilsages looks like you got it..

OpenStudy (anonymous):

so the slope would be 0, but how do I find the slope of the line that is parallel and perpendicular to the equation?

OpenStudy (texaschic101):

put the equation in y = mx + b form... 2x - 2y = - 6 -2y = - 6 - 2x y = (-2x - 6)/-2 y = x + 3 For a parallel line, the slope is the same....so the slope for this is 1. Do you understand ?

OpenStudy (texaschic101):

if there is no number in front of x, it is 1.....always. The same with y, if there is no number in front of it, it is 1.

OpenStudy (anonymous):

ok got it

OpenStudy (texaschic101):

so the slope for parallel line is the same as the slope in the original equation, which is 1

OpenStudy (anonymous):

but how do I find the line that's perpendicular

OpenStudy (luigi0210):

perpendicular slopes are the negative reciprocals

OpenStudy (texaschic101):

now, the perpendicular line. The slope for the perpendicular line is the negative reciprocal.

OpenStudy (anonymous):

what does that mean, negative opposites?

OpenStudy (texaschic101):

what is the negative reciprocal of 1 ?

OpenStudy (anonymous):

-1

OpenStudy (texaschic101):

example.....2.....negative reciprocal is - 1/2

OpenStudy (luigi0210):

\[x=-\frac{ 1 }{ x }\] This is a negative reciprocal

OpenStudy (anonymous):

o

OpenStudy (anonymous):

so it would be 1=-1/1

OpenStudy (anonymous):

the negative reciprocal would be -1

OpenStudy (texaschic101):

-1/1.......rise -1, run + 1.....@Luigi0210 ...would it just be -1 or -1/1

OpenStudy (anonymous):

what?

OpenStudy (texaschic101):

@Luigi0210 ........

OpenStudy (luigi0210):

it's -1

OpenStudy (texaschic101):

thats what I thought....just making sure.....thank you Luigi...your a good person

OpenStudy (luigi0210):

Haha, thank you, you are too :P

OpenStudy (anonymous):

wait so am i using the formula x=-1/x to find all the perpendicular slopes?

OpenStudy (texaschic101):

you are using the negative reciprocal........for instance -3.....parallel would be -3, perpendicular would be 1/3.

OpenStudy (texaschic101):

-3 is the same as -3/1....so flip it and make it positive

OpenStudy (anonymous):

so if it were 3 i'd flip it 1/3 and make it negative, -1/3?

OpenStudy (anonymous):

so for all of them I just flip it and make it positive?

OpenStudy (texaschic101):

3/1...negative reciprocal is - 1/3.. No....only make it positive if it is negative.....if it is positive, make it negative....make it the opposite

OpenStudy (anonymous):

ok so flip it and make it opposite?

OpenStudy (texaschic101):

let me show you... this is just for perpendicular... slope 4.......negative reciprocal is - 1/4 slope -5.....negative reciprocal is 1/5 flip it and change the sign

OpenStudy (texaschic101):

yes....flip it and make it opposite

OpenStudy (anonymous):

got it thanks so much!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

slope intercept form is y=mx+b right?

OpenStudy (texaschic101):

your welcome...remember parallel is the same and perpendicular, you flip it and make it opposite. You first have to put it in y = mx + b form. Got it ?

OpenStudy (anonymous):

yes got it :)

OpenStudy (anonymous):

ty you r awesome!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (texaschic101):

if you have any more troubles with it, just flag me or Luigi, he is good. We will be happy to help :)

OpenStudy (anonymous):

Ok thank you so much will do!!!!!!!!!!!!!!

OpenStudy (luigi0210):

well I'm not good.. only she is :P

OpenStudy (texaschic101):

yes you are.....

OpenStudy (anonymous):

you both rock @Luigi0210 @texaschic101

OpenStudy (anonymous):

ty so much!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (luigi0210):